# Benchmark Performance of Homomorphic Polynomial Public Key Cryptography for Key Encapsulation and Digital Signature Schemes

### 05 Jan 2024 v1

@@ 4 Security Summary
On the other side, a comprehensive security analysis of HPPK DS can be found in our paper [5]. The optimal attack involves key recovery, specifically finding moduli $S_{1},S_{2}$ of the homomorphic encryption key, with a complexity of $\mathcal{O}(S_{1}S_{2}/p)=\mathcal{O}(2^{2L}/p)$. Once $S_{1}$ and $S_{2}$ are known, determining other elements of the HPPK private key becomes relatively straightforward by combining the HPPK DS public key with intercepted genuine signatures. The intriguing factor is the Barrett parameter $R=2^{k}$: how big is the $k$? A larger $k$ results in a larger public key size. Table 2 illustrates four configurations for each security level, three for $m=2$ with $K=L+32,L+64,2L$, respectively, and the last configuration for $m=1$ with $K=L+32$.
The table demonstrates that different configurations do not affect private key and signature sizes, only influencing public key sizes. For $m=2$, as the Barrett parameter $k$ decreases from $K=2L$ to $K=L+32$, the public key size significantly decreases from 544B to 376B for level I, from 792B to 528B for level III, and from 1040B to 680B for level V, respectively. On the other hand, the public key size is reduced by almost $50\%$ when $m$ is changed from 2 to 1, as shown in the 4th configuration for all three security levels.
@@
Table 2: The key and signature sizes in bytes, as provided by the HPPK DS scheme for the proposed parameter sets, are determined based on the optimal complexity of $\mathcal{O}(2^{2L}/p)$. In this context, we choose the Barrett parameter $R$ to be $32$ bits longer than $S_{1}/S_{2}$, and the hidden ring size is set to be $L=2\times|p|_{2}+16$. All data is presented in bytes. The configuration is defined as $(n,\lambda,m,L,log_{2}R).$

### 10 Jan 2024 v2

@@ 4 Security Summary
Within the HPPK DS framework, Kuang et al. introduced a key recovery attack leveraging specific public key values, denoted as $s_{1}$ and $s_{2}$ [5]. The initial method for determining moduli $S_{1}$ and $S_{2}$ exhibited a computational complexity of $\mathcal{O}(S_{1}S_{2}/p)=\mathcal{O}(2^{2L}/p)$. This paper presents an optimized key recovery attack directly utilizing public key elements $\mu_{ij}$ and $\nu_{ij}$, defined as:
$\mu_{ij}=\lfloor\frac{R\cdot p_{ij}}{S_{1}}\rfloor\longrightarrow p_{ij}=\lceil\frac{S_{1}\cdot\mu_{ij}}{R}\rceil$ (10)
$\nu_{ij}=\lfloor\frac{R\cdot q_{ij}}{S_{2}}\rfloor\longrightarrow q_{ij}=\lceil\frac{S_{2}\cdot\nu_{ij}}{R}\rceil$
This streamlined attack involves iteratively searching for $S_{1}$ within the range $2^{L-1}$ to $2^{L}$, calculating $p_{ij}$ using the known public key $\mu_{ij}$, and recalculating $\mu_{ij}$ with $S_{1}$ and $p_{ij}$. If the recomputed $\mu_{ij}$ matches the public key $\mu_{ij}$, the attacker deterministically identifies the private values $S_{1}$ and $p_{ij}$. The computational complexity of this approach is $\mathcal{O}(2^{L-1})$. A similar procedure applies to $S_{2}$ and $q_{ij}$, resulting in a total complexity of $\mathcal{O}(2^{L})$.
Additionally, the remaining private key elements can be effortlessly determined by intercepting genuine signatures with the known values of $p_{ij}$, $q_{ij}$, $S_{1}$, and $S_{2}$. This optimized key recovery attack exhibits a complexity of $\mathcal{O}(2^{L})$, showcasing significantly enhanced computational efficiency compared to the previous approach with a complexity of $\mathcal{O}(2^{2L}/p)$.
Table 2 demonstrates that different configurations do not affect private key and signature sizes, only influencing public key sizes. For $m=2$, as the Barrett parameter $k$ decreases from $K=2L$ to $K=L+32$, the public key size significantly decreases from 544B to 376B for level I, from 792B to 528B for level III, and from 1040B to 680B for level V, respectively. On the other hand, the public key size is reduced by almost $50\%$ when $m$ is changed from 2 to 1, as shown in the 4th configuration for all three security levels.
@@
Table 2: The key and signature sizes in bytes, as provided by the HPPK DS scheme for the proposed parameter sets, are determined based on the optimal complexity of $\mathcal{O}(2^{L})$. In this context, we choose the Barrett parameter $R$ to be $32$ bits longer than $S_{1}/S_{2}$, and the hidden ring size is set to be $L=2\times|p|_{2}+16$. All data is presented in bytes. The configuration is defined as $(n,\lambda,m,L,log_{2}R).$